OpenStudy (anonymous):

HELP HELP HELP PLEASE!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!! make 'a' the subject

OpenStudy (anonymous):

\[R _{1}=R _{2}(1+\alpha \theta)\]

OpenStudy (anonymous):

Steps please :) :)

OpenStudy (anonymous):

steps 1 divide both sides by R_2 2 subtract 1 from both sides 3 divide by \[\theta\]

OpenStudy (anonymous):

does r1 divded by r2 equal r1 or r1/r2? @Jonask

OpenStudy (anonymous):

when you divide by \(R_2\) both sides you get \(\dfrac{R_1}{R_2}\) on the left

OpenStudy (anonymous):

can i have help with step 3? i have \[\frac{ R _{1}-1 }{ R _{2} } = \alpha \theta \]

OpenStudy (anonymous):

how do i divide by the 0? @Jonask

OpenStudy (anonymous):

step 2 isnt very good,lets work on that first ,when u subtract both sides by 1 ,u dont subtract from the top part alone ut the whole fraction eg. subtract 1 from \(\dfrac{3}{2}\) means \(\dfrac{3}{2}-1 \) and not \(\dfrac{3-1}{2}\)

OpenStudy (anonymous):

oh i see

OpenStudy (anonymous):

can u show me your step 2

OpenStudy (anonymous):

subtract 1 from \(\dfrac{R_1}{R_2}\)

OpenStudy (anonymous):

\[\frac{ R _{1} }{ R _{2} }-1 = \alpha \theta \]

OpenStudy (anonymous):

is that right?

OpenStudy (anonymous):

yes,thats right,now you can actually make it simper,by using this rule \[\dfrac{a}{b}-1=\frac{a-b}{b}\]

OpenStudy (anonymous):

does that mean it would be .. \[\frac{ R _{1} }{ R _{2} }- 1=\frac{ R _{1}-R _{2} }{ R _{2}}\]

OpenStudy (anonymous):

yes yes,finally we need to divide by \(\theta\),this is not zero(0),its called theta \[\dfrac{R_1-R_2}{R_2}=\alpha\theta\] divide both sides by \(\theta\)

OpenStudy (anonymous):

okay i get this all but how do you divide a fraction?

OpenStudy (kc_kennylau):

(a/b)/c = a/(bc)

OpenStudy (anonymous):

yes use the rule above ^^ by kc_kennylau

OpenStudy (anonymous):

d\[\frac{ R _{1}- R _{2} }{ R _{2}\theta }\]

OpenStudy (anonymous):

Is that what you mean?

OpenStudy (anonymous):

yes exactly,you are done ,and correct

OpenStudy (anonymous):

oh yay! thamnk you @Jonask so sos so so much! wow helped so much. It's so hard doing homework without the teacher. But you helped me so much! and thanks @kc_kennylau for the equation thats going to be so useful

OpenStudy (kc_kennylau):

Thanks for the acknowledgement and no problem at all :)

OpenStudy (anonymous):

no problem @muppz ,your welcome